Fama’s Llamas has a WACC of 10.4 percent. The company’s cost of equity is 13.4 percent, and its pretax cost of debt is 8.8 percent. The tax rate is 38 percent. What is the company’s target debt–equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1
Weighted average cost of capital (WACC) = [(S/S+B)*Rs + (B/S+B)*Rb(1-tc)]
S = equity, B = debt, Rs = Cost of equity, Rb = cost of debt,
tc = corporations tax rate
WACC = .104
Rs = .134
Rb = .088
tc = .38
Let (S/S+B) = x
1 - (S/S+B) = (B/S+B) = 1-x
B/S = (1-x)/x
.104 = x*.134 + (1-x)*.088*(1-.38)
.104 = x*.134 + (1-x)*.088*(.62)
.104 = .134x + (1-x)*(.05456)
.104 = .134x + .05456 - .05456x
0.04944 = .07944x
x = .04944/.07944
x = .622356
1- x = 1 - .622356
1 - x = .377644
B/S = .377644/.622356
B/S = .606796
The company's target DEBT/EQUITY ratio is .6068.